OpenStudy (anonymous):

Find a nonzero vector u with initial point P(-1, 3, -5) such that u has the same direction as v = (4, 2, 5). Suppose that the first component of u is u1 = 8

OpenStudy (anonymous):

This is BS. There's no actual significance of the initial point.

OpenStudy (anonymous):

Basically, a vector \(\langle a,b,c\rangle \) and a vector \(\langle na, nb, nc\rangle\) have the same direction. It's just that the second vector's magnitude is \(n\) times the first. You will have to use that fact here.

OpenStudy (anonymous):

You are given that \(\langle 4,2,5\rangle \) and \(\langle 8, x_0 , y_0\rangle \) have the same direction.

OpenStudy (anonymous):

Do you see how the vector \(\langle 2\times 4 , 2 \times 2, 2 \times 5\rangle \) has the same direction as \(\langle 4,2,5\rangle\)?

OpenStudy (anonymous):

ok i get it you use u=kv, since u=8 you can now solve k which is 8,4,10? :)

OpenStudy (anonymous):

Precisely!

OpenStudy (anonymous):

thank you :D